# LongVILA: Scaling Long-Context Visual Language Models for Long Videos

### v2

@@ Abstract
Long-context capability is critical for multi-modal foundation models. We introduce LongVILA, a full-stack solution for long-context vision-language models, including system, model training, and dataset development. On the system side, we introduce the first long-context Multi-Modal Sequence Parallelism (MM-SP) system that enables long training and inference, enabling 2M context length training on 256 GPUs without any gradient checkpointing. MM-SP is 2.1×2.1$\times$2.1 × - 5.7×5.7$\times$5.7 × faster than ring sequence parallelism and $\times1$ 1.4×1.4$\times$1.4 × faster than Megatron context parallelism + tensor parallelism in text-only settings. Moreover, it seamlessly integrates with Hugging Face Transformers. For model training, we propose a five-stage pipeline comprising alignment, pre-training, short supervised fine-tuning, context extension, and long supervised fine-tuning. On datasets, we construct large-scale visual language pre-training datasets and long video instruction-following datasets to support our multi-stage training process. LongVILA extends the number of frames of VILA from 8 to 1024, and improves the long video captioning score from 2.00 to 3.26 (1.6x), achieving 99.5% accuracy in 1400-frames video (274k context length) needle-in-a-haystack. LongVILA-8B demonstrates consistent accuracy improvements on long videos in the VideoMME benchmark as the number of frames increases.
@@ 1 Introduction
A full-stack design is crucial for long-context Vision-Language Models (VLMs). Training large models is typically a complex, systematic endeavor that demands both data engineering (Betker et al., 2023; Ouyang et al., 2022; Zhou et al., 2024) and system-software co-design (Lepikhin et al., 2020; Chowdhery et al., 2023; Shoeybi et al., 2019; Brown et al., 2020; Dehghani et al., 2023). Unlike text-only LLMs, VLMs (e.g., LLaVA (Liu et al., 2023b)) often require distinct model architectures and flexible distributed training strategies. Additionally, long-context modeling necessitates not only long-context data to fully utilize the model’s capabilities (Fu et al., 2024b; Chen et al., 2023) but also infrastructure capable of supporting memory-intensive long-context training (Li et al., 2021; Jacobs et al., 2023; Li et al., 2023a). Therefore, a meticulously planned full-stack design, encompassing system, data, and pipeline, is indispensable for long-context VLMs.
@@
In this work, we introduce LongVILA, a comprehensive solution for long-context VLMs. For training infrastructure, we establish an efficient and user-friendly framework, namely Multi-Modal Sequence Parallelism (MM-SP), which supports training memory-intensive long-context VLMs. For training pipeline, we implement a five-stage training curriculum as Figure 1: (1) multi-modal alignment, (2) large-scale pre-training, (3) short supervised fine-tuning, (4) context extension for LLMs, and (5) long supervised fine-tuning. For inference, MM-SP addresses the challenge of KV cache memory usage, which becomes a bottleneck with very long sequences, by deploying an inference mode within our MM-SP framework to support long-context multi-modal language deployment. For training data, we focus on two key requirements: (1) the collection of a large volume of high-quality visual-language data for large-scale pre-training, and (2) the acquisition of high-quality long-context visual-language instruction-following data to unlock long-range modeling capabilities.
By efficiently increasing the number of video frames using LongVILA, the performance on VideoMME (Fu et al., 2024a) and long video captioning tasks consistently improves (Figure 2). The LongVILA model, trained on 1024 frames, achieves 99.5% accuracy in the needle-in-a-haystack experiments with 1400 frames, equivalent to a context length of 274k tokens. Our MM-SP system can efficiently scale the context length up to 2 million tokens without gradient checkpointing, achieving 2.1$\times$ to 5.7$\times$ speedup compared to ring sequence parallelism, and 1.1$\times$ to 1.4$\times$ compared to Megatron context parallelism + tensor parallelism.
@@ 2 Background and Related work
Long-context training examples often exceed the memory capacity of a single device. To address this issue, the sequence parallelism paradigm has been widely adopted in the text-only LLM community, distributing a single sequence across multiple devices. Specifically, Ring-style systems Li et al. (2021; 2023a); Liu et al. (2023a) use Point-to-Point (P2P) communication primitives to collectively compute the attention module, while DeepSpeed-Ulysses Jacobs et al. (2023) employs an All-to-All (A2A) primitive to alternate between sharding the sequence dimension and the attention head dimension during attention computation. Our work is the first to design and implement a sequence parallelism system for visual language models. Moreover, we have thoroughly examined both A2A and P2P primitives and developed a novel hybrid sequence parallelism strategy to optimize training throughput. Concurrent work, such as (Gu et al., 2024), also explores integrating these two SP strategies. However, it is limited to conventional LLM training and does not extend to multi-modal scenarios due to the lack of support for complex attention masks and variable-length input sequences.
@@ 3 Multi-Modal Sequence Parallelism
Training long-context Vision-Language Models (VLMs) results in substantial memory demands. For example, in our Stage 5 long video training (Figure 1), a single sequence comprises around 200K tokens for 1024 video frames, which exceeds the memory capacity of a single GPU. The most widely used open-source solution, fully sharded data parallelism, does not distribute the activations generated by a single sequence, making it unsuitable for our needs. Consequently, we developed a custom system based on sequence parallelism(Li et al., 2021; 2023a; Liu et al., 2023a; Jacobs et al., 2023), a technique commonly employed in existing foundation model systems to optimize text-only LLM training. However, we discovered that existing systems are neither efficient nor scalable enough to handle our long-context VLM workloads.
Figure 3: MM-SP incorporates our innovative 2D attention technique, which utilizes intra-node All-to-All (All2All) and inter-node Point-to-Point (P2P) communication to transfer keys and values (KV). It also includes a novel sharding strategy that balances the computational load between the image encoder and the language modeling stages. In contrast, Ring-style SP (Liu et al., 2023a; Li et al., 2021; 2023a) relies on P2P communication for both intra-node and inter-node settings, leading to significant communication overhead. Additionally, these systems are designed for text-only modalities, without optimization for the workload of an image encoder. Intra-node and inter-node bandwidths were measured in internal H100 clusters.
@@ 3.2 Multi-Modal Sequence Parallelism Training Mode
After identifying the limitations in existing systems, we conclude that an ideal multi-modal sequence parallelism approach should prioritize efficiency and scalability by addressing both modality and network heterogeneity, and should also be capable of scaling beyond the number of attention heads.
Figure 4: Workflow of Multi-Modal Sequence Parallelism with a Batch Size (bs) of 4 and a Sequence Parallel Size (SP_Size) of 3. To facilitate sequence parallelism for multi-modal inputs, we developed a customized sharding strategy that ensures balanced workload distribution and compatibility with our 2D-attention mechanism.
@@
Figure 5: Comparison of RingAttention, ZigZag-RingAttn, and 2D-Attention. The blue block indicates communication between QKV, while the black frame represents local attention computation within each SP group rank (SP_Size = 3).
2D attention parallelism. To address network heterogeneity and achieve scalability, we combine the strengths of ring-style sequence parallelism (SP) and Ulysses SP. Specifically, we consider parallelization across the sequence dimension or the attention head dimension as ”1D SP.” Our method extends this by parallelizing computation across both the attention head and sequence dimensions, transforming a 1D SP process group into a 2D mesh comprising independent Ring (P2P) and Ulysses (A2A) process groups. For instance, as illustrated on the left in Figure 3, to enable an 8-degree sequence parallelism across 2 nodes, we construct a 4$\times$2 communication mesh using 2D-SP. In this setup, the A2A process group, with a size of 4, distributes the QKV tensors according to the head dimension and re-partitions them along the sequence dimension within each node. Simultaneously, the P2P process group, with a size of 2, transfers the partitioned KV chunks between nodes.
Additionally, to further explain how ZigZag-RingAttnbalances computation and how the 2D-attention mechanism operates, we depict the attention computation schedule using different methods in Figure 5. In this example, the sequence length is 8 and the global SP degree is 4. Due to the triangular structure of causal attention computations, RingAttention experiences a computation imbalance, where rank 0 becomes idle after the first round while rank 3 continues computing through all stages. ZigZag-RingAttnaddresses this by reordering input sequence tokens along the sequence dimension to achieve load balance. Our 2D-attention mechanism uses a ring parallel degree of 2 and a head parallel degree of 2, resulting in an effective global sequence parallel degree of 4. This approach also incorporates a workload balancing strategy within the ring-based process group and uses the All-to-All operation to distribute QKV tensors across devices based on the head dimension, ensuring efficient and balanced computation.
@@ 3.3 Multi-Modal Sequence Parallelism Inference Mode
4 LongVILA Training Data and Pipeline
4.1 Overall Training Pipeline
As shown in Figure 1, in our pipeline, there are five training stages, i.e., Stage 1: multi-modal alignment, Stage 2: large-scale pre-training, Stage 3: supervised fine-tuning, Stage 4: context extension for LLM, Stage 5: long supervised fine-tuning. Stage 1, 2, and 3 are following VILA Lin et al. (2023b), to firstly bridge the gap between LLM and vision encoder, and then pre-training on larger datasets. In Stage 1, only the multi-modal projector is trainable with others frozen. In Stage 2, we frozen the vision encoder and training LLM and the multi-modal projector. In Stage 3, we fully fine-tune the model for short data instruction following, e.g., image and short video datasets. Afterwards, we extend the context length of LLM with text-only dataset in a continued pre-training manner in Stage 4. In Stage 5, we enhance the instruction following abilities by long video supervised fine-tuning. It is noted that all parameters are trainable in the final stage.
@@ 4.2 Stage1&2&3: Alignment, Pre-training, and Short Supervised Fine-tuning
4.3 Stage4: Context Extension for LLMs
4.4 Stage5: Long Supervised Fine-tuning
Long video instruction following dataset To facilitate the fine-tuning of long videos, we also constructed a new, dedicated dataset for long video training, consisting of 15,292 videos for each one. We use the original long videos from the Shot2Story dataset (Han et al., 2023). Each video includes different questions and answers: one for generating captions and another for answering questions, enabling diverse applications in video understanding. Figure 7 illustrates the process for generating instruction-following datasets from long videos. Initially, the long video is segmented into shorter clips, each approximately 10 seconds in duration. These clips are then independently annotated with descriptive captions utilizing the VILA-1.5 model. Subsequently, an LLM is employed to generate question-and-answer pairs derived from the captions of these clips. The generated questions encompass summarization and other queries relevant to the comprehensive understanding of long video content.
5 Evaluation
5.1 Training and inference system
@@ 5.1.2 Inference system
We evaluated our inference system against HuggingFace Pipeline parallelism using a single node with 8 H100 GPUs and the 8B model (Table  3). Our system achieves an 8.2×8.2$\times$8.2 × speedup compared to HuggingFace Pipeline on 8xH100 GPUs. This significant improvement is primarily due to HuggingFace Pipeline inference activating only one GPU at a time, whereas our method leverages all GPUs to compute jointly. Table 3 compares the maximum supported sequence length, where our method supports sequences that are 2.9×2.9$\times$2.9 × longer than those supported by HuggingFace Pipeline. Specifically, during 96K sequence length inference, the maximum sequence length supported, HuggingFace Pipeline stores 80GB of activations on the first GPU and only 18GB on the remaining GPUs. This imbalanced allocation of activations limits the maximum sequence length that can be processed.
@@ 6 Conclusion
We introduce LongVILA, a comprehensive full-stack solution for long-context visual language models, encompassing distributed systems, model training, and data engineering. Our system efficiently scales context length up to 2 million tokens, achieving speedups of 2.1$\times$ to 5.7$\times$ compared to ring sequence parallelism and 1.1$\times$ to 1.4$\times$ compared to Megatron context parallelism + tensor parallelism. Additionally, we have meticulously curated long video datasets and implemented a five-stage training pipeline. Our LongVILA-8B model extends the feasible frame count from 8 to 1024, precisely capturing fine-grained information from 2-hour needle-in-a-haystack videos and delivering promising results in both video QA and captioning tasks.

### v3

@@ Abstract
Long-context capability is critical for multi-modal foundation models, especially for long video understanding. We introduce LongVILA, a full-stack solution for long-context visual-language models by co-designing the algorithm and system. For model training, we upgrade existing VLMs to support long video understanding by incorporating two additional stages, i.e., long context extension and long supervised fine-tuning. However, training on long video is computationally and memory intensive. We introduce the long-context Multi-Modal Sequence Parallelism (MM-SP) system that efficiently parallelizes long video training and inference, enabling 2M context length training on 256 GPUs without any gradient checkpointing. LongVILA efficiently extends the number of video frames of VILA from 8 to 1024, improving the long video captioning score from 2.00 to 3.26 (out of 5), achieving 99.5% accuracy in 1400-frame (274k context length) video needle-in-a-haystack. LongVILA-8B demonstrates consistent accuracy improvements on long videos in the VideoMME benchmark as the number of frames increases. Besides, MM-SP is 2.1×2.1$\times$2.1 × - 5.7×5.7$\times$5.7 × faster than ring sequence parallelism and $\times1$ 1.4×1.4$\times$1.4 × faster than Megatron with context parallelism + tensor parallelism. Moreover, it seamlessly integrates with Hugging Face Transformers.
@@ 1 Introduction
A full-stack design is crucial for long-context Vision-Language Models (VLMs). Training large models is typically a complex, systematic endeavor that demands both data engineering (Betker et al., 2023; Ouyang et al., 2022; Zhou et al., 2024) and system-software co-design (Lepikhin et al., 2020; Chowdhery et al., 2023; Shoeybi et al., 2019; Brown et al., 2020; Dehghani et al., 2023). Unlike text-only LLMs, VLMs (e.g., LLaVA (Liu et al., 2023b)) often require distinct model architectures and flexible distributed training strategies. Additionally, long-context modeling necessitates not only long-context data to fully utilize the model’s capabilities (Fu et al., 2024b; Chen et al., 2023) but also infrastructure capable of supporting memory-intensive long-context training (Li et al., 2021; Jacobs et al., 2023; Li et al., 2023a). Therefore, a full-stack design, encompassing training pipeline and system, is indispensable for long-context VLMs.
@@
In this work, we introduce LongVILA, a comprehensive solution for long-context VLMs. For training pipeline, we implement a five-stage training curriculum as Figure 1: (1) multi-modal alignment, (2) large-scale pre-training, (3) short supervised fine-tuning, (4) context extension for LLMs, and (5) long supervised fine-tuning. For training system, we establish an efficient and user-friendly framework, namely Multi-Modal Sequence Parallelism (MM-SP), which supports training memory-intensive long-context VLMs. For inference, MM-SP addresses the challenge of KV cache memory usage, which becomes a bottleneck with very long sequences, by deploying an inference mode within our MM-SP framework to support long-context multi-modal language deployment.
By efficiently increasing the number of video frames using LongVILA, the performance on VideoMME (Fu et al., 2024a) and long video captioning tasks consistently improves (Figure 2). The LongVILA model, trained on 1024 frames, achieves 99.5% accuracy in the needle-in-a-haystack experiments with 1400 frames, equivalent to a context length of 274k tokens. Our MM-SP system can efficiently scale the context length up to 2 million tokens without gradient checkpointing, achieving 2.1$\times$ to 5.7$\times$ speedup compared to ring sequence parallelism, and 1.1$\times$ to 1.4$\times$ compared to Megatron with context parallelism + tensor parallelism.
@@ 2 Background and Related work
Long-context training examples often exceed the memory capacity of a single device. To address this issue, the sequence parallelism paradigm has been widely adopted in the text-only LLM community, distributing a single sequence across multiple devices. Specifically, Ring-style systems Li et al. (2021; 2023a); Liu et al. (2023a) use Point-to-Point (P2P) communication primitives to collectively compute the attention module, while DeepSpeed-Ulysses Jacobs et al. (2023) employs an All-to-All (A2A) primitive to alternate between sharding the sequence dimension and the attention head dimension during attention computation. Ulysses generally achieves higher throughput than Ring-style SP due to its more efficient A2A communication primitive and larger, unsegmented computation blocks. However, its scalability is limited by the number of attention heads. Recently, USP (Fang & Zhao, 2024) was introduced as the first to integrate Ring-style SP and Ulysses SP, combining the strengths of both approaches. LoongTrain (Gu et al., 2024) further optimizes communication and placement strategies to enhance training efficiency. Following (Fang & Zhao, 2024; Gu et al., 2024), we extend the system to multi-modal scenarios to accommodate complex attention masks and variable-length input sequences. Our work is the first to design and implement a sequence parallelism system for visual language models.
3 LongVILA Training Pipeline
As shown in Figure 1, in our pipeline, there are five training stages, i.e., Stage 1: multi-modal alignment, Stage 2: large-scale pre-training, Stage 3: supervised fine-tuning, Stage 4: context extension for LLM, Stage 5: long supervised fine-tuning. Stage 1, 2, and 3 are following VILA Lin et al. (2023b), to firstly bridge the gap between LLM and vision encoder, and then pre-training on larger datasets. In Stage 1, only the multi-modal projector is trainable with others frozen. In Stage 2, we frozen the vision encoder and training LLM and the multi-modal projector. In Stage 3, we fully fine-tune the model for short data instruction following, e.g., image and short video datasets. Afterwards, we extend the context length of LLM with text-only dataset in a continued pre-training manner in Stage 4. In Stage 5, we adopt our MM-SP system (§4) to enhance the instruction following abilities by long video supervised fine-tuning. It is noted that all parameters are trainable in the final stage.
3.2 Stage4: Context Extension for LLMs
3.3 Stage5: Long Supervised Fine-tuning
Long video instruction following To facilitate the fine-tuning of long videos, we also constructed a new, dedicated dataset for long video training, consisting of 15,292 videos for each one. We use the original long videos from the Shot2Story dataset (Han et al., 2023). Each video includes different questions and answers: one for generating captions and another for answering questions, enabling diverse applications in video understanding. Figure 3 illustrates the process for generating instruction-following datasets from long videos. Initially, the long video is segmented into shorter clips, each approximately 10 seconds in duration. These clips are then independently annotated with descriptive captions utilizing the VILA-1.5 model. Subsequently, an LLM is employed to generate question-and-answer pairs derived from the captions of these clips. The generated questions encompass summarization and other queries relevant to the comprehensive understanding of long video content.
Once we acquired the long video dataset, applying it for supervised fine-tuning introduced new challenges, primarily due to the substantial number of frames in each sample—often ranging in the hundreds or even thousands. For instance, a single sequence from 1400 video frames can encompass around 274k tokens. Existing data-parallel training systems struggle to handles such extensive contexts. To address this limitation, we developed the MM-SP system (§4) to efficiently train long-context visual language models.
@@ 4 Multi-Modal Sequence Parallelism
Training long-context Vision-Language Models (VLMs) results in substantial memory demands. The most widely used open-source solution, fully sharded data parallelism, does not distribute the activations generated by a single sequence, making it unsuitable for our needs. Consequently, we developed a custom system based on sequence parallelism(Li et al., 2021; 2023a; Liu et al., 2023a; Jacobs et al., 2023), a technique commonly employed in existing foundation model systems to optimize text-only LLM training. However, we discovered that existing systems are neither efficient nor scalable enough to handle our long-context VLM workloads.
Figure 5: Sharding strategy and communication pattern of MM-SP. For sharding strategy, Ring SP is designed for text-only modalities, without optimization for the workload of an image encoder. Our MM-SP implements a novel sharding strategy that balances the computational load between the image encoder and the language modeling stages. For communication pattern, Ring SP (Liu et al., 2023a; Li et al., 2023a) (top) relies on P2P communication for both intra-node and inter-node settings, resulting in underutilization of intra-node bandwidth. MM-SP (bottom) adopts 2D-Attention (Fang & Zhao, 2024; Gu et al., 2024) mechenism which utilizes intra-node All-to-All (All2All) and inter-node Point-to-Point (P2P) communication to transfer keys and values (KV), enhancing the efficiency of intra-node NVLink utilization. The bandwidth is for H100.
@@ 4.2 Multi-Modal Sequence Parallelism Training Mode
After identifying the limitations in existing systems, we conclude that an ideal multi-modal sequence parallelism approach should prioritize efficiency and scalability by addressing both modality and network heterogeneity, and should also be capable of scaling beyond the number of attention heads. To achieve this, we adopt 2D-attention (Fang & Zhao, 2024; Gu et al., 2024) mechanism for sequence parallelism. For instance, as illustrated on the left in Figure 5, to enable an 8-degree sequence parallelism across 2 nodes, we construct a 4$\times$2 communication mesh using 2D-SP. In this setup, the A2A process group, with a size of 4, distributes the QKV tensors according to the head dimension and re-partitions them along the sequence dimension within each node. Simultaneously, the P2P process group, with a size of 2, transfers the partitioned KV chunks between nodes. Additionally, to further explain how the 2D-attention mechanism operates, we depict the attention computation schedule using different methods in Figure 11.
Figure 6: Workflow of Multi-Modal Sequence Parallelism with a Batch Size (bs) of 4 and a Sequence Parallel Size (SP_Size) of 3. To facilitate sequence parallelism for multi-modal inputs, we developed a customized sharding strategy that ensures balanced workload distribution and compatibility with SP communucation.
@@ 4.3 Multi-Modal Sequence Parallelism Inference Mode
5 Evaluation
5.1 Training and inference system
@@ 5.1.2 Inference system
We evaluated our inference system against HuggingFace Pipeline parallelism using a single node with 8 H100 GPUs and the 8B model (Table  3 and Figure 9 (b)). Our system achieves an 8.2×8.2$\times$8.2 × speedup compared to HuggingFace Pipeline on 8xH100 GPUs. This significant improvement is primarily due to HuggingFace Pipeline inference activating only one GPU at a time, whereas our method leverages all GPUs to compute jointly. Table 3 compares the maximum supported sequence length, where our method supports sequences that are 2.9×2.9$\times$2.9 × longer than those supported by HuggingFace Pipeline. Specifically, during 96K sequence length inference, the maximum sequence length supported, HuggingFace Pipeline stores 80GB of activations on the first GPU and only 18GB on the remaining GPUs. This imbalanced allocation of activations limits the maximum sequence length that can be processed.
@@ 6 Conclusion
We introduce LongVILA, a comprehensive full-stack solution for long-context visual language models, encompassing model training pipeline and distributed system. Based on our curated long video datasets and five-stage training pipeline, our LongVILA-8B model extends the feasible frame count from 8 to 1024, precisely capturing fine-grained information from 2-hour needle-in-a-haystack videos and delivering promising results in both video QA and captioning tasks. Additionally, our system efficiently scales context length up to 2 million tokens, achieving speedups of 2.1$\times$ to 5.7$\times$ compared to ring sequence parallelism and 1.1$\times$ to 1.4$\times$ compared to Megatron context parallelism + tensor parallelism.
@@ Appendix A Appendix
Figure 11: Comparison of RingAttention  Liu et al. (2023a), ZigZag-RingAttn Zhu (2023), and 2D-Attention (Fang & Zhao, 2024). The blue block indicates communication between QKV, while the black frame represents local attention computation within each SP group rank. The sequence length is 8 and the global SP degree is 4. Due to the triangular structure of causal attention computations, RingAttention experiences a computation imbalance, where rank 0 becomes idle after the first round while rank 3 continues computing through all stages. ZigZag-RingAttn addresses this by reordering input sequence tokens along the sequence dimension to achieve load balance. The 2D-Attention mechanism uses a ring parallel degree of 2 and a head parallel degree of 2, resulting in an effective global sequence parallel degree of 4. This approach also incorporates a workload balancing strategy within the ring-based process group and uses the All-to-All operation to distribute QKV tensors across devices based on the head dimension, ensuring efficient and balanced computation.